OpenStudy (anonymous):

If z^2=x^2+y^2 with z>0, dx/dt=3, and dy/dt=6, find dz/dt when x=4 and y=3.

OpenStudy (anonymous):

@ranga @zepdrix

OpenStudy (anonymous):

@phi

OpenStudy (phi):

take the derivative with respect to t (implicit differentiation) do you know how to do that ?

OpenStudy (anonymous):

yes but thats what i thought i should do but then whats the point of all the other info

OpenStudy (phi):

First, what do you get for the d/dt of that equation ?

OpenStudy (anonymous):

there is no t in the equation so how do i do d/dt

OpenStudy (phi):

use the "rule" d/dt of a variable (x , y or z) is dx/dt, dy/dt or dz/dt use the chain rule where needed: example d/dt x^2 = 2 x dx/dt

OpenStudy (anonymous):

ok. so its 2zz'=2xx'+2yy'

OpenStudy (phi):

yes. notice that with the original equation z^2 = x^2 + y^2 you can find z (given x and y) they tell you x' and y' so fill in the numbers in your derivative and solve for z'

OpenStudy (anonymous):

how do i find z and y

OpenStudy (phi):

x=4 and y=3.

OpenStudy (phi):

z^2=x^2+y^2

OpenStudy (phi):

*** that should answer your question " whats the point of all the other info?" ****

OpenStudy (anonymous):

z'=30/sqrt21?

OpenStudy (phi):

x=4, y= 3, z=5, x'=3, y'=6 5 z'=4*3+3*6=12+18= 30 z'= 6

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

you should remember that 3^2 + 4^2= 5^2 (3,4,5 right triangle triplet) not sure how you get 21 out of that

OpenStudy (anonymous):

yeah i really messed up somewhere. thanks though